10. Given the following two measurements of the equilibrium constant for a reaction, calculate AH for the reaction. T.°C 30.0 50.0 5.2 x 10 15.4 x 10
Answer #1

we have:

T1 = 30 oC

=(30+273)K

= 303 K

T2 = 50 oC

=(50+273)K

= 323 K

K1 = 5.2*10^3

K2 = 1.54*10^5

we have below equation to be used:

ln(K2/K1) = (Ea/R)*(1/T1 - 1/T2)

ln(1.54*10^5/5.2*10^3) = ( Ea/8.314)*(1/303.0 - 1/323.0)

3.3883 = (Ea/8.314)*(2.044*10^-4)

Ea = 137850 J/mol

Ea = 137.8 KJ/mol

Answer: 137.8 KJ/mol

Feel free to comment below if you have any doubts or if this answer do not work. I will edit it if you let me know